# Apparate: Rethinking Early Exits to Tame Latency-Throughput Tensions in ML Serving

### v1

@@ 3.3 Latency-Focused Ramp Adjustments
To cope with this uncertainty, our guiding intuition is that, subject to the same accuracy constraint, later ramps almost always exhibit higher exit rates than earlier ones. [^2] For a single input, being able to exit at an early ramp does not guarantee exit capabilities at later ramps [34]. However, later-ramp exit rates were always higher than earlier ones for our workloads, and we note that Apparate only uses this property for search efficiency, not correctness. The reason is that late ramps have the luxury of leveraging more of an original model’s computations when making a prediction. Importantly, this implies that a candidate ramp’s exit rate is bound by the exit rate of the closest downstream ramp.
@@
Figure 12: Computing upper-bound exit rates for candidate ramps. Blue dots show previously active ramps (+/+/-+ / - indicates positive/negative utility), while orange dots show candidates.
Until now, we have only discussed how Apparate handles scenarios with at least one negative ramp utility. In the event that all ramps exhibit positive utilities, Apparate enters a low-risk probing phase to determine if latency savings can grow by injecting (or shifting to) earlier ramps. There are two scenarios for this. If ramp budget remains, we add a ramp immediately before the existing ramp with highest utility (while keeping that ramp to preserve its exiting wins). In contrast, if no ramp budget remains, we shift the ramp with the lowest utility score one position earlier, leaving the most positive ramp untouched. Additions are incorporated into the configuration using the same process as above.
4 Implementation
Apparate is implemented as a layer atop TensorFlow-Serving [39] and Clockwork [22] (using PyTorch [7]) and includes the components described in 3 written as Python modules in $\sim$7500 lines of code. Although we chose these platforms for our current implementation, we note that Apparate is not limited to them and its techniques can be implemented in any inference platform. Importantly, Apparate entirely leverages the scheduling and queuing mechanisms of the underlying framework. Original models are ingested in the ONNX format [6] and compiled for performance. Ramp training (during bootstrapping) uses the first 10% of each dataset following a 1:9 split for training and validation; the remaining 90% of each dataset is used for evaluation.
5 Evaluation
@@
    Apparate lowers 25th percentile and median latencies by 40.5-91.5% and 70.2-94.2% for CV, and 16.0-37.3% and 10.0-24.2% for NLP workloads, compared to original (non-EE) models.
@@
    Unlike existing EE models that unacceptably worsen accuracies and tail latencies by up to 23.9% and 11.0%, Apparate consistently meets specified accuracy and tail latency constraints.
@@ 5.1 Methodology
Models. We consider 10 models (across 4 families) that cover popular architectures and a variety of model sizes in both CV and NLP. For CV, we use the ResNet{18, 50, 101} residual models, as well VGG{11, 13, 16} models that follow a chained (linear) design. All of these models are pre-trained on ImageNet and from the PyTorch Model Zoo [43]; we further fine-tune the models to our video scenes using a random sampling of 10% of frames across our dataset. For NLP, we consider 3 encoder-only transformers from the BERT family – BERT-base, BERT-large, and Distilbert [45] (a variant of BERT-base that was compressed via distillation) – as well as a decoder-only large language model: GPT2-medium. These models span 66-345 million parameters, were collected from HuggingFace [30], and were pre-trained (without more fine-tuning) on Yelp reviews [8] that are separate from any evaluation dataset we use.
@@
Workloads. CV workloads comprise real-time object classification (people, cars) on 8 one-hour videos used in recent video analytics literature [11, 27]. The videos were sampled at 30 frames per second, and collectively span day/night from diverse urban scenes; for each, we perform classification.
@@
NLP workloads focus on sentiment analysis using two datasets: Amazon product reviews [10] and IMDB movie reviews [41]. To the best of our knowledge, there do not exist public streaming workloads for NLP classification, so we convert these datasets into streaming workloads as follows. For Amazon, we follow the order of product categories in the original dataset, but within each category, we keep reviews only from frequent users (i.e., those with $>$1k reviews) and order streaming by user (250k requests in total). For IMDB, we follow the order of reviews in the original dataset, but stream each in sentence by sentence (180k requests in total). We then define arrival patterns for these ordered requests by using the Microsoft Azure Functions (MAF) as in prior serving work [22]. Specifically, to ensure meaningful and realistic workloads despite the high degree of variation in runtime across our models, we paired each model with a randomly selected trace snippet from the set that met the following criteria: (1) number of requests match that in our largest dataset, and (2) queries per second should not result in $>$20% dropped requests with vanilla serving for the given model and selected SLO (described next).
@@
Parameter configurations. Given our focus on interactivity, we cope with heterogeneity in model runtimes by setting SLOs to be 2$\times$ each model’s inference time with batch size 1 in our main experiments. This results in SLOs between 13-204 ms, which match the ranges considered in prior work [49, 40, 22]; Table 5 in §A lists the specific SLO values per model, and we study the effect of SLO on Apparate in §5.4. Unless otherwise noted, results use 1% for Apparate’s accuracy constraint (in line with industry reports [42]) and a ramp budget of 2% impact on worst-case latency; we consider other values for both parameters in §5.4.
@@
Setup. Experiments were conducted on a dedicated server with one NVIDIA RTX A6000 housing 48GB of memory, one AMD EPYC 7543P 32-Core CPU, and 256GB DDR4 RAM. We run experiments with two serving platforms: TensorFlow-Serving [39] and Clockwork [22]. We primarily present results with Clockwork due to space constraints, but note that report trends hold for both platforms; we compare cross-platform results in §5.4.
@@
Metrics and baselines. Our main metrics of evaluation are classification accuracy and per-request latencies. Accuracy is defined as the percentage of inputs that are assigned the correct label as defined by each original (i.e., non-EE) model. Per-request latency is measured as the time between when a request arrives at a serving platform, and when its response is released by the platform. We mainly compare Apparate with two baselines: (1) original models without EEs (vanilla) running in serving platforms, and (2) optimal EEs as defined in §2.2, i.e., assuming all inputs exit at their earliest possible ramps with non-exiting inputs incurring no ramp overheads. We compare Apparate to existing EE strategies in §5.3.
@@
Figure 13: Median latency savings compared to vanilla models. Bars show median workload with error bars for min-max.
@@
Figure 14: Evaluating Apparate’s impact on tail latency (running with 2% budget) compared to vanilla serving. Bars show median workload, with error bars spanning min-max.
@@
Figure 15: Apparate (with 2% budget) vs. vanilla models on NLP workloads. “-V” indicates serving using the vanilla model. Note that “-V” curves per plot mostly overlap since they use the same timing trace and no exiting; minor discrepancies are only due to the varying number of inputs across workloads.
@@
Figure 16: Apparate vs. optimal exiting on NLP workloads with the Amazon dataset.
5.2 Overall Results
@@
Figures 1316 compare Apparate with vanilla model serving and optimal exiting across our workloads. Overall, Apparate significantly lowers latencies compared to serving vanilla models, while always adhering to the imposed 1% accuracy constraint. For instance, median speedups range from 40.5-91.5% (2.7-30.5 ms) for CV workloads, and jump to 70.2-94.2% (5.2-31.4 ms) at the 25th percentiles. NLP workloads follow a similar pattern, with median and 25th percentile savings ranging from 10.0-24.2% (3.9-25.3 ms) and 16.0-37.3% (4.8-53.2 ms), respectively. Importantly, across all workloads, Apparate’s tail latency (and thus impact on throughput) always falls within its granted budget (2% here) and is most often negligible (Figure 14).
@@
Beyond this, there are several important trends to note. First, Apparate’s raw latency savings grow with increased model sizes, e.g., 25th percentile wins of 53.2ms, 28.4ms, 14.3ms, 5.5ms for GPT-2, BERT-large, BERT-base, and Distilbert-base on the Amazon dataset. This is because only the results exit models with Apparate, with inputs running to completion; thus, latency savings pertain entirely to serving times (not queuing delays), with exit impact being higher as model size (and thus, runtime) grows. Relative (%) latency savings follow the same pattern for CV workloads, e.g., Apparate’s median wins grow by 13.8% and 5.3% moving from the smallest to the biggest models in the ResNet and VGG families. However, relative wins remain relatively stable in NLP models, e.g., 15.8% and 13.7% for GPT2 and BERT-large on Amazon Reviews. The reason for this difference is in the effectiveness of the models in each domain. Results and task performance are largely similar across the considered CV models, enabling Apparate to inject ramps early in (even larger) models. In contrast, results and task efficacy are far better with the larger models in NLP; thus, Apparate’s ramps fall in similar (relative) positions across the models.
@@
Second, Apparate’s wins are larger for CV workloads than NLP workloads for two reasons. As previously noted, CV workloads use lighter models and lower request rates (bound by video fps), and thus incur far lower queuing delays. More importantly, in contrast to CV where spatiotemporal similarities across frames (and thus, requests) are high due to physical constraints of object motion in a scene, NLP requests exhibit less continuity, e.g., back-to-back reviews are not constrained in semantic similarity. The effects on Apparate’s adaptation of EE configurations are that (1) past data is less representative of future data, and (2) the duration until subsequent adaptation is shorter.
Lastly, Apparate’s wins for NLP are consistently higher with the Amazon dataset versus the IMDB dataset. The reason is that the IMDB dataset exhibits ‘harder’ inputs that require later-ramp exits. For instance, for GPT-2, average exit locations for the offline optimal exit strategy are at layers 2.2 and 2.7 for Amazon and IMDB, respectively; these numbers are 2.1 and 2.5 for BERT-base.
(a) CV
(b) NLP
Figure 17: Impact of SLOs on Apparate’s wins.
@@
Comparisons with optimal. As shown in Figure 13, latency savings with Apparate for CV workloads largely mirror those of the optimal that tunes exiting decisions based on perfect knowledge of the upcoming workload, e.g., median savings are within 20.5% of the optimal. In contrast, building on the discussion above, the limited continuity across inputs in NLP workloads leads to a wider gap of 73.1-83.2% at the median (Figure 16). To further characterize Apparate’s performance on these workloads, we also consider a more realistic online optimal algorithm that relaxes the following elements. First, rather than per-sample adaptation of thresholds and ramps, ramp adjustments are set to operate only as fast as model definitions in the GPU can be updated. Second, rather than using perfect knowledge of upcoming inputs, decisions are made using only recent (historical) data; for each threshold or ramp adaptation decision, we tune based on the past {20, 40, 80} batches of inputs and select the decision that performs best on the upcoming data. As shown in Figure 16, Apparate’s median latency savings are within 16.9-52.1% of this (more) realistic optimal exiting strategy.
@@
Varying SLOs. SLOs affect serving system batching and queuing decisions, and thus Apparate’s benefits. To study this effect, we considered SLOs for each model that were 2$\times$ and 4$\times$ those in our default experiments (Table 5). Generally, higher SLOs induce larger serving batch sizes and higher per-request queuing delays (as inputs wait for the previous, slower inference batch to complete); this dampens Apparate’s relative latency savings which target model runtimes, but not queuing delays. Figure 17 illustrates this effect. For instance, median latency savings for GPT-2 drop from 16.3% to 6.8% as SLO grows by 4$\times$. Note that, to illustrate this trend for CV workloads, we upsampled each video to 120 fps. The reason is that our serving platforms are work-conserving and, at 30 fps, they are able to consistently schedule jobs with batch size 1 and low queuing delays given the low runtimes of the considered CV models.
5.3 Comparison with Existing EE Strategies

### v2

@@ 3.3. Latency-Focused Ramp Adjustments
To cope with this uncertainty, our guiding intuition is that, subject to the same accuracy constraint, later ramps almost always exhibit higher exit rates than earlier ones. The reason is that late ramps have the luxury of leveraging more of an original model’s computations when making a prediction. Importantly, this implies that a candidate ramp’s exit rate is bound by the exit rate of the closest downstream ramp; note that this is not a formal guarantee (Kaya et al., 2019), and Apparate uses this for search efficiency (not correctness).
@@
Figure 11. Computing upper-bound exit rates for candidate ramps. Blue dots show previously active ramps (+/+/-+ / - indicates positive/negative utility), while orange dots show candidates.
Until now, we have only discussed how Apparate handles scenarios with at least one negative ramp utility. In the event that all ramps exhibit positive utilities, Apparate enters a low-risk probing phase to determine if latency savings can grow by using earlier ramps. If ramp budget remains, we add a ramp immediately before the existing ramp with highest utility (while keeping that ramp to preserve its exiting wins). If not, we shift the ramp with the lowest utility score one position earlier, leaving the most positive ramp untouched.
3.4. Supporting Generative Large Language Models
Recent efforts have incorporated EEs into generative language models (e.g., CALM (Schuster et al., 2022) and FREE (Bae et al., 2023) for T5 (Raffel et al., 2020), LayerSkip (Elhoushi et al., 2024) for Llama (Touvron et al., 2023)) to enhance their response times in interactive applications. While the general challenges with EEs persist – lack of runtime adaptation for time-varying accuracy and latency savings – the auto-regressive nature of generative models poses a unique challenge for Apparate. Unlike with classification where inputs can be processed to completion independently, each token generated by an LLM depends on the preceding ones in the sequence, requiring their full key-value (KV) states. Thus, when a token $T1$ exits using a ramp, the generation delay begins to accumulate for the next token, $T2$; yet, $T2$’s forward pass through the layers after the ramp cannot begin until $T1$ passes through the same layers to generate KV states. The effect is potential harm on time-per-token (TPT) distributions.
To regain latency savings from EEs, Apparate draws inspiration from recent parallel decoding techniques (Bae et al., 2023; Leviathan et al., 2023; Elhoushi et al., 2024). As a token exits from a ramp, Apparate does not immediately compute the remaining model layers, and instead only accumulates its hidden states at that ramp. The remaining computations are executed in parallel alongside the first non-exiting token that is encountered thereafter at that ramp. For concreteness, consider a scenario with 2 tokens, $T1$ and $T2$. Assume $T1$ can exit at ramp $R1$, while $T2$ must proceed later in the model. $T1$ will exit at $R1$, and $T2$ will immediately begin processing. However, once $T2$ fails to exit at $R1$, the remaining layers for $T1$ are run alongside (i.e., batched with) $T2$’s remaining layers. Taken together, $T1$ achieves per-token latency savings from exiting, while $T2$ incurs a very mild (§4.3) penalty from the higher batch size.
In addition to improved TPT latencies and compute efficiency (due to batching effects on GPUs), parallel decoding grants Apparate token-level feedback for exiting decisions relative to the original (non-EE) model. Specifically, for each parallel decoding instance, Apparate collects per-token feedback up until the first token whose exit result deviates from the original model; feedback for subsequent tokens is discarded as it may reflect cascading errors from inter-token dependencies. This feedback guides Apparate’s threshold and ramp tuning strategies from §3.23.3.
4. Evaluation
@@
    Apparate lowers 25th percentile and median classification latencies by 40.5–91.5% and 70.2–94.2% for CV, and 16.0–37.3% and 10.0–24.2% for NLP workloads, compared to original (non-EE) models. These wins are 5.7–66.6% larger than two-layer inference systems using compressed models for ‘easy’ inputs (Wang et al., 2023; Canel et al., 2019). Median time-per-token wins are 22.6–77.9% for generative scenarios.
@@
    Unlike existing EE models that unacceptably worsen accuracies and tail latencies by up to 23.9% and 11.0% for classification, Apparate consistently meets accuracy and tail latency constraints. This carries to generative scenarios: Apparate’s tail latency is 1.8–2.4% lower than existing EE models which lower accuracies up to 5.5%.
@@ 4.1. Methodology
Models. For classification, we consider 10 models (across 4 families) that cover popular architectures and diverse sizes. For CV, we use the ResNet{18, 50, 101} residual models, as well VGG{11, 13, 16} models that follow a chained (linear) design. All of these models are pre-trained on ImageNet and from the PyTorch Model Zoo (PyTorch, 2023). For NLP, we consider 3 encoder-only transformers from the BERT family – BERT-base, BERT-large, and Distilbert (Sanh et al., 2019) (a variant of BERT-base that was compressed via distillation) – as well as a decoder-only LLM: GPT2-medium. These models span 66–345 million parameters, were collected from HuggingFace (HuggingFace, 2023), and were pre-trained on Yelp reviews (yel, 2024). We also consider quantized versions of these BERT models in §4.2. For generative scenarios, we use the T5-large (Raffel et al., 2020) encoder-decoder LLM (from prior EE work (Schuster et al., 2022; Bae et al., 2023)) with 770 million parameters and the Llama2 decoder-only LLMs with 7 and 13 billion parameters, both pretrained from HuggingFace.
@@
Workloads. CV workloads comprise real-time object classification (people, cars) on 8 one-hour videos from recent video analytics literature (Agarwal and Netravali, 2023; Hsieh et al., 2018). The videos were sampled at 30 frames per second, and span day/night from urban scenes.
@@
NLP classification workloads focus on sentiment analysis using two datasets: Amazon product reviews (ama, 2013) and IMDB movie reviews (Pal et al., 2020). To the best of our knowledge, there do not exist public streaming workloads for NLP classification, so we convert these datasets into streaming workloads as follows. For Amazon, we follow the order of product categories in the original dataset, but within each category, we keep reviews only from frequent users (i.e., those with $>$1k reviews) and order streaming by user (250k requests in total). For IMDB, we follow the order of reviews in the original dataset, but stream each in sentence by sentence (180k requests in total). We then define arrival patterns using the Microsoft Azure Functions (MAF) as in prior work (Gujarati et al., 2020; Li et al., 2023). To cope with the large variation in runtime across our models, we paired each model with a randomly selected trace snippet from the set that met the following criteria: (1) number of requests match that in our largest dataset, and (2) queries per second should not result in $>$20% dropped requests with vanilla serving for the given model and selected SLO (described below).
Our generative workloads use two datasets: CNN/DailyMail (See et al., 2017) for text summarization and SQuAD (Rajpurkar et al., 2016) for question answering. As in prior work (Schuster et al., 2022; Bae et al., 2023), we assign request arrival times that follow a Poisson distribution, configured to saturate our computing resources.
@@
Parameter configurations. Given our focus on interactivity, we cope with heterogeneity in model runtimes by setting SLOs for classification to be 2$\times$ each model’s inference time with batch size 1 in our main experiments. This results in SLOs between $\sim$10–200 ms, which match the ranges used in prior work (Shen et al., 2019; Padmanabhan et al., 2022; Gujarati et al., 2020); Table 5 in §7 lists the specific SLO values, and we study the effect of SLO on Apparate in §4.5. Unless otherwise noted, results use 1% for Apparate’s accuracy constraint and a ramp budget of 2% impact on worst-case latency; we consider other parameter values in §4.5.
@@
Setup. Experiments were conducted on dedicated servers with NVIDIA RTX A6000 GPUs housing 48GB of memory, AMD EPYC 7543P 32-Core CPUs, and 256GB DDR4 RAM. We run experiments with two serving platforms for classification: TensorFlow-Serving (Olston et al., 2017) and Clockwork (Gujarati et al., 2020). For space, we primarily present results with Clockwork, but note that reported trends hold for both platforms; we compare cross-platform results in §4.5. For generative workloads, we use the HuggingFace Pipelines inference engine (hf_, 2024).
@@
Metrics and baselines. For classification, our main metrics are accuracy (of inputs assigned correct label as per the non-EE model) and per-request response latency (including queuing). For generative models, sequence accuracy is measured using ROUGE-L (text summarization) and F1 (question answering) scores. Latency is measured using time-per-token (TPT) distributions. We compare Apparate with the following baselines: (1) original models without EEs (vanilla), (2) optimal EEs as defined in §2.2, i.e., assuming all inputs exit at their earliest possible ramps, with no ramp overheads, (3) two-layer inference systems (Canel et al., 2019; Wang et al., 2023) that invoke full models only when compressed ones cannot generate high-confidence outputs, and (4) existing, non-adaptive EE strategies (§4.4).
@@
Figure 12. Median latency savings vs. vanilla models. Bars show median savings across all workloads; error bars are min-max.
@@
Figure 13. Evaluating Apparate’s impact on tail latency (running with 2% budget) vs. vanilla serving. Bars show the median savings across all workloads with error bars spanning min-max.
@@
Figure 14. Apparate (with 2% budget) vs. vanilla models (“-V”) on NLP classification workloads. “-V” curves per plot mostly overlap since they use the same timing trace and no exiting; minor discrepancies are due to the varying number of inputs across workloads.
@@
Figure 15. Apparate vs. optimal exiting on NLP classification workloads with the Amazon dataset.
4.2. Results for CV and NLP Classification
@@
Figures 1215 compare Apparate with vanilla model serving and optimal exiting across our workloads. Overall, Apparate significantly lowers latencies compared to serving vanilla models, while always adhering to the imposed 1% accuracy constraint. For instance, median speedups range from 40.5–91.5% (2.7–30.5 ms) for CV workloads, and jump to 70.2–94.2% (5.2–31.4 ms) at the 25th percentiles. NLP workloads follow a similar pattern, with median and 25th percentile savings ranging from 10.0–24.2% (3.9–25.3 ms) and 16.0–37.3% (4.8–53.2 ms). Importantly, across all workloads, Apparate’s tail latency (and thus impact on throughput) always falls under its 2% budget, and is most often negligible (Figure 13).
@@
Beyond this, there are several important trends to note. First, Apparate’s raw latency savings grow with increased model sizes, e.g., 25th percentile wins of 53.2ms, 28.4ms, 14.3ms, 5.5ms for GPT-2, BERT-large, BERT-base, and Distilbert-base on the Amazon Reviews. This is because only results (not inputs) exit models with Apparate, so latency savings pertain entirely to serving times (not queuing delays) and grow as model size/runtime grows. Relative (%) latency savings follow the same pattern for CV workloads, e.g., Apparate’s median wins grow by 13.8% and 5.3% moving from the smallest to the biggest models in the ResNet and VGG families. However, relative wins remain relatively stable in NLP models, e.g., 15.8% and 13.7% for GPT2 and BERT-large on Amazon Reviews. The difference is due to the effectiveness of the models in each domain. Results and task performance are largely similar across the CV models, enabling Apparate to inject ramps early in (even larger) models. In contrast, results are far better with the larger models in NLP; thus, Apparate’s ramps fall in similar (relative) positions across the models.
@@
Second, Apparate’s wins are larger for CV workloads than NLP workloads for two reasons. As previously noted, CV workloads use lighter models and lower request rates (bound by video fps), and thus incur far lower queuing delays. More importantly, in contrast to CV where spatiotemporal similarities across frames (and thus, requests) are high due to physical constraints of object motion in a scene, NLP requests exhibit less continuity, e.g., back-to-back reviews are not constrained in semantic similarity. The effects on Apparate’s adaptations are that (1) past data is less representative of future data, and (2) the duration until subsequent adaptation is shorter.
Other compute optimizations. To further illustrate Apparate’s ability to run alongside existing compute efficiency optimizations – a key goal of its design (§3) – we ran experiments using post-training Int8 quantized Bert-base and Bert-large. Overall, we observe that Apparate’s speedups largely persist, which median and 25th percentile wins of 7.3–19.4% and 6.9–31.1%. The mild dip in speedups relative to non-quantized Bert models (Figure 14) is a result of quantization’s reduction in model overparameterization, which early exiting aims to capitalize on for select inputs.
(a) CV: VGG11, VGG13.
(b) NLP: Distilbert, Bert-base
Figure 16. Apparate (“-A”) vs. two-layer inference systems (“-T”).
Figure 17. Impact of SLOs on Apparate’s wins.
@@
Comparisons with optimal. As shown in Figure 12, latency savings with Apparate for CV workloads largely mirror those of the optimal that tunes exiting decisions based on perfect knowledge of the upcoming workload, e.g., median savings are within 20.5% of the optimal. In contrast, the limited continuity across inputs in NLP workloads leads to a wider gap of 65.4–78.5% at the median (Figure 15). To further characterize Apparate’s performance on these workloads, we also consider a more realistic online optimal algorithm that relaxes the following elements. First, rather than per-sample adaptation of thresholds and ramps, ramp adjustments are set to operate only as fast as model definitions in the GPU can be updated. Second, rather than using perfect knowledge of upcoming inputs, decisions are made using only recent (historical) data; we tune based on the past {20, 40, 80} batches of inputs and select the one that performs best on the upcoming data. As shown in Figure 15, Apparate’s median latency savings are within 4.9–17.2% of this (more) realistic optimal strategy.
Comparisons with two-layer inference. We compare Apparate with in-house implementations of Tabi (Wang et al., 2023) (NLP) and FilterForward (Canel et al., 2019) (CV), which employ compressed models on all inputs, and pass only those inputs with low confidence results to the base model. Acceptable confidence scores from compressed models are configured to ensure both systems operate within the same accuracy loss budget as Apparate. Note that our evaluation is favorable to these systems: we ignore overheads from hosting the compressed models, compute overheads for data pruning (e.g., word pruning in Tabi), and queuing for batch formation between compressed and base models. As shown in Figure 16, across 3 representative workloads, Apparate delivers 5.7–66.6% and 20.9–42.0% lower median and P95 latencies compared to these baselines. For easy inputs, Apparate delivers latency wins by enabling ramps early enough (i.e., within the first third) in the base model to run faster than the baselines’ compressed models. For hard inputs, tail latencies with Apparate are capped by its 2% budget; in contrast, the baselines add the entire compressed model runtime per input.
(a) T5
(b) Llama2
Figure 18. Comparison of Apparate against T5 generative LLM and FREE (Bae et al., 2023) (left), and against Llama generative LLM and its optimal variant (right). Bars represent medians with 25-95th percentiles.
@@
Varying SLOs. We considered SLOs for each model that were 2$\times$ and 4$\times$ those in our default experiments (Table 5). Generally, higher SLOs induce larger serving batch sizes and higher per-request queuing delays; this dampens Apparate’s relative latency savings (Figure 17) which target model runtimes, but not queuing delays. For instance, median latency savings for GPT-2 drop from 13.2% to 6.8% as SLO grows by 4$\times$. Note that, to illustrate this trend for CV workloads, we upsampled each video to 120 fps. The reason is that our serving platforms are work-conserving and, at 30 fps, they are able to consistently schedule jobs with batch size 1 and low queuing delays given the low model runtimes.
4.3. Results for Generative Scenarios
Figure 18 compares Apparate with the vanilla T5-large LLM across our workloads. As shown, Apparate lowers median and 25th percentile TPT by 70.4–77.9% and 74.6–78.0%, respectively. However, at the 95th percentile, Apparate incurs 2.6–8.4% higher TPT due to ramp overheads and (more so) added delay from parallel decoding (non-exiting tokens run alongside the remaining computation for previously-exited tokens). Unlike with classification, formally bounding tail overheads online is hard because variable sequence lengths make TPT values unpredictable in vanilla generative serving. We also evaluated Apparate on the larger Llama models for question answering. Figure 18 shows that Apparate reduces the median and 25th percentile TPT by 22.6–37.4% and 25.4–40.3%. Notably, the latency wins increase as model size grows.
Comparison to optimal. Similar to classification, we define the optimal EE strategy as exiting each token at the earliest possible ramp that generates the correct value (ignoring delays for generating the remaining KV states). Apparate’s median TPT for T5 and Llama is within 9.0–16.2% and 2.1%–7.7% of the optimal, respectively. The smaller gap relative to NLP classification stems from two elements. First, accuracy here is measured at the sequence level, granting more flexibility for exiting decisions at individual tokens. Further, auto-regressive generation (with shared state across tokens) grants more continuity than between requests in NLP classification, boosting adaptation benefits.
4.4. Comparison with Existing EE Strategies